Session: May 1997 (session 2)
Frequency: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Tue Jun 10 02:03:55 METDST 1997) Success
Sometimes antenna offsource for few sec. from 160-17:15. Problems to air-conditioner at the end of experiment.

Effelsberg (dgraham at Tue Jun 10 10:09:49 METDST 1997) Minor failures
scan lost at 11:15 , tape lost vacuum

Torun (K.Borkowski at Tue Jun 10 11:52:00 METDST 1997) Minor failures
12:22 to 13:45 off source (observations stopped due to local thunderstorm)

Jodrell (Alastair Gunn at Mon Jun 16 11:26:10 METDST 1997) Success
No known problems.

Onsala (John Conway at Wed Jun 18 13:49:44 METDST 1997) Success
No known problems

Medicina (Fredrik T Rantakyro at Mon Jul 21 11:19:55 METDST 1997) Minor failures
The ANTAB file for the experiment can be found on VLBEER in the MAY97 directory with the name EF3B.ANTAB. The file is ready to be used with the AIPS tasks ANTAB/APCAL. Read the ANTAB file for any specific comments regarding this experiment.

MPIfR (Scott Aaron at Mon Aug 3 14:39:33 METDST 1998) Minor failures
Correlation completed. Noto data correlated twice due to a clock offset error in the correlator control file. Standard tracking problems with the Torun data. Jodrell data only on forward tape passes.